OpenStudy (anonymous):

Jeeves, the valet, was promised a salary of $8000 and a car for a year of service. When Jeeves left the job after 7 months of service, he received the car and $1600 as his prorated salary. What was the dollar value of the car?

OpenStudy (anonymous):

Please help! This is the only one on my quiz I'm struggling on...

OpenStudy (hoblos):

value of the car = x (8000 +x)/12 = (1600+x)/7 solve it and you will get x=7360

OpenStudy (anonymous):

Okay. Thanks so much!!

OpenStudy (hoblos):

MY PLEASURE XD